OpenStudy (anonymous):

HELP?!!?! What is the equation in point-slope form of the line passing through (-2, -5) and (2, 3)?

OpenStudy (dangerousjesse):

The slope is 2 and y-intecept is -1. Can you make that an equation?

OpenStudy (anonymous):

i have no idea how to do this i honestly just need the answer.

OpenStudy (dangerousjesse):

Sadly I can't just give you the answer.

OpenStudy (anonymous):

ok

OpenStudy (dangerousjesse):

To find the point-slope equation, you simply find the y-intercept, which, in this case, is negative one, and the slope (2). Since slope is \(\frac{y_2-y_1}{x_2-x_1}=\textsf{slope}\), you'll use the inverse operation of \(\frac{y}{x}=\textsf{slope}\), which is \(x\times \textsf{slope}\)\(=y\), or \(2x\) and you'll add the y-intercept to that to find y, or \(y=2x-1\)